# sections/MRM-supplementary

### v1

@@ Abstract
Results. WALINET is significantly faster and needs 8s for high-resolution whole-brain MRSI, compared to 42 minutes for conventional HLSVD+L2. Quantitative analysis shows WALINET has better performance than HLSVD+L2: 1) more lipid removal with 41% lower NRMSE, 2) better metabolite signal preservation with 71% lower NRMSE in simulated data, 155% higher SNR and 50% lower CRLB in in-vivo data. Metabolic maps obtained by WALINET in healthy subjects and patients show better gray/white-matter contrast with more visible structural details.
@@ 1 Introduction
Proton magnetic resonance spectroscopic imaging (1H-MRSI) has great potential as a metabolic imaging technique that can measure the intrinsic metabolite concentrations across the whole-brain without the need to administer molecular agents 1. Proton-based MR spectroscopy, compared to X-nuclei spectroscopy, has a significantly higher SNR and is ubiquitously available on all MRI scanners equipped with standard hardware and software. Consequently, it represents the vast majority of MRS performed in clinical studies and research applications. In specific pathology it allows the detection of metabolic abnormalities in the brain before anatomical lesions are visible 2. Therefore, it is valuable for disease investigation such as brain tumor identification, classification and treatment response assessment 3. It further enables the study of neurochemistry alterations and disease mechanisms in neuropsychiatry 4. However, the clinical potential of 1H-MRSI is not fully realized due to complex experimental factors that reduce data quality such as large artifacts from the overwhelming water and lipid signals 5, 6. Hence, efficient solutions to these technical problems can have great and widespread impact for the use of 1H-MRSI in clinical applications.
@@
The removal of lipid signal originating from the scalp region presents with even greater challenges 7. The lipid spectrum is complicated, having multiple peaks that fully overlap with aliphatic region and are 1-2 orders of magnitude stronger than metabolite peaks. Furthermore, the lipid signal originates from scalp areas with very inhomogeneous $B_{0}$ field, hence they are much broader than water and metabolite peaks from inside the brain.
Methods enabling suppression of water and lipid contamination can be categorized into several groups 7: techniques leveraging specific RF-pulse or sequence designs to invert, nullify or saturate water or lipid resonances 8, 9, 10, 11, 12, 13, 14, 15 approaches utilizing dedicated hardware to spoil the scalp signal 16, 17, and post-acquisition methods employing spatial or spectral priors for lipid contamination removal. The latter category in the case of lipids removal can be divided in methods using lipid signal extrapolation 18, dual-density reconstruction 19, 20, lipid-basis penalty 21, 22, subspace reconstruction methods based on specific spatial supports and spectral decomposition 23, 24, 25, while in the case of water removal includes Hankel matrix singular value decomposition 26, subspace reconstruction 23, Lowner tensorization 27, and water-basis penalty 28.
@@
Recently, closed form solutions have been developed for lipid removal 29, 30. Closed form solutions are derived by applying a linear operator on each individual spectrum and are faster compared to iterative lipid suppression algorithms. However, the linear operator method has some drawbacks too: 1) it is subject specific, 2) requires accurate anatomical mask delineation, and 3) may not work if the orthogonal assumption and linear superposition between lipids and metabolites is not met..
@@
In recent years, deep learning-based methods applied to mitigate MRS challenges have gained popularity, enabling robust applications and eliminating the need for complex parameter optimizations. Thereby, applications can be categorized into deep learning based artifact removal, denoising, lowrank 31, 32, 33 and spectral quantification 34, 35, 36, 37, 38.
@@ 2 Methods
The problem can be formulated using two distinct inputs $x_{1}$ and $x_{2}$: 1) the original MRS spectrum ($x_{1}$) containing metabolite signal $m$ contaminated with lipid $l$ and water $w$ signal, and 2) the spectrum ($x_{2}$) subjected to a projection onto the lipid subspace, $\mathds{1}-\mathcal{L}$, using lipid L2 regularization 29 approach with $\mathcal{L}=(\mathds{1}+\beta\mathbf{LL^{H}})^{-1}$. $\mathbf{L}$ is a matrix containing the in-vivo lipid signal obtained from the scalp mask of each subject and $\beta$ is the regularization parameter. The derived operator $\mathcal{L}$ is an approximation of a projection whose kernel is given by the linear span of $\mathbf{L}$. $\mathds{1}-\mathcal{L}$, with $\mathds{1}$ being an identity matrix, is a projection onto the lipid subspace span($\mathbf{L}$) (see references for further details). A separate lipid projection operator is calculated for each subject. The network inputs are defined as
@@
WALINET employs a Y-Net convolutional neural network structure 39, depicted in Figure 1, and characterized by the use of two encoders instead of one. Note, that the same network architecture can be used and trained only for lipid removal, which we call LIPNET (LIpid neural NETwork).
The Y-Net architecture enables enhanced contextual understanding by integrating features from different branches, leading to improved understanding of complex structures within the data. The results show an improvement compared to the U-Net 39 and enable the encoding of different features in each encoder 40.
@@
Each encoder ($\mathcal{E}$) and decoder ($\mathcal{D}$) of the Y-Net $(\mathcal{Y}$) comprised four convolutional blocks, each consisting of two convolutional layers, PReLU activation functions 41, dropout with a rate of 0.01, and MaxPooling/upsampling with a factor of 2. Skip connections are implemented between the encoders and decoder. An additional convolutional block is incorporated in the bottleneck region and after the decoder. The outputs of the encoders are concatenated and forwarded to the bottleneck convolutional block in the decoder.
@@
The training spectra were generated through a multi-step process, where metabolite spectra were simulated and combined with experimentally measured lipid and water spectra from in-vivo human data. This approach was chosen because: 1) metabolite spectra can be realistically simulated for an extremely large range of very diverse parameters, 2) the lipid and water spectra are affected by a complex combination of experimental factors which are hard to be fully accounted in simulations, hence more realistic spectra can be extracted from measured data. First, spectra for 25 common 1H metabolites were simulated using a physical model for the coupled spin systems42, 43. Based on these modes, an extensive dataset of $10^{6}$ metabolite spectra was generated by simulation with a parameter distribution that encompassed a wide range of concentrations, linewidths, noise levels, and baselines. The metabolite concentrations were distributed according to a normal distribution with a standard deviation five times greater than the mean of 1 (arbitrary units), with truncation to zero to ensure positive values. The simulation included also variations in frequency offset (-150 to 150 Hz), Voigt linewidth (4 to 50 Hz), and signal-to-noise ratio (SNR from 1 to 10). Random baselines were introduced, characterized by 10 broad gaussian components.
Second, we extracted a collection of lipid spectra from voxels within the scalp region obtained from in-vivo MRSI datasets. Representative head/brain/scalp masks and B0 field maps are presented in Supplementary Figure 1. To augment the in-vivo lipid distribution, the amplitude of lipid signal from the lipid mask was varied with a scaling factor, spanning a broad interval from $1^{-2}$ to $1^{3}$ to match the lipid contamination inside the brain.
Lastly, to create the final training input spectra ($x_{1}$) for WALINET the simulated metabolite spectra ($m$) and experimentally-derived lipid & water spectra ($x_{2}$) were randomly combined. Note, that in the case of LIPNET the input training spectra combined simulated metabolite and experimentally-derived lipid spectra, without the water signal.
The training dataset was generated using water and lipid signals extracted from 19 subjects, including 2 glioma patients. The MRSI data were acquired with the 3D 1H-FID-ECCENTRIC sequence described in paragraph 2). $10^{5}$ lipid & water spectra were extracted from each subject, resulting in $1.9\times 10^{6}$ total lipid & water spectra used for training. Additionally, a validation dataset included 4 other subjects.
@@
The network was trained for 400 epochs, using the Adam 44 optimizer with a learning rate of 0.01, which was quartered every 50 epochs. The exponential decay rates of the first and second momentum of Adam $\beta_{1}$ $\beta_{2}$ were set to 0.9 and 0.999. The network was implemented using PyTorch 2.0.1 and CUDA 11.7 packages in Python 3.8.13. The model training was performed on a PowerEdge R7525 server (Dell) with 64 CPU cores (AMD EPYC7542 2.90GHz, 128M Cache, DDR4-3200), 512 GB CPU RAM (RDIMM, 3200MT/s), 3 GPU NVIDIA Ampere A40 (PCIe, 48GB GPU RAM) running Rocky Linux release 8.8 (Green Obsidian).
@@
In-vivo MRSI data were acquired with 2D 1H-FID Cartesian phase encoded 45 and 3D 1H-FID ECCENTRIC 46 pulse sequences using a 7T MR scanner (MAGNETOM Terra, Siemens Healthineers, Forchheim, Germany) and a 1Tx/32Rx head coil (NovaMedical, Wilmington, MA, USA).
@@
3D 1H-FID-ECCENTRIC 46 was acquired on 21 subjects with 64x64x31 matrix, 220x220x105 mm3 FoV, 3.4x3.4x3.4 mm3 voxel size, spectral bandwidth of 2326 Hz, and 453 FID points.
@@
For comparison of WALINET and LIPNET performance, conventional nuisance signal removal was performed with HLSVD for water 26 and L2 regularization for lipids. 29. The HLSVD retained the 32 largest eigenvalues of the Hankel matrix and the water removal was applied in the frequency range of 4.7ppm±$\pm$0.5ppm. L2 regularization used lipid signals extracted from the skull mask of the subject with the regularization parameter $\beta$ individually adjusted for each subject to achieve a mean absolute diagonal value of its lipid suppression operator $\mathcal{L}=(\mathds{1}+\beta\mathbf{LL^{H}})^{-1}$ at an arbitrary value of 0.938,
@@
This value was selected as the optimal trade-off between minimizing metabolite alteration and maximizing lipid suppression, Following the water and lipid removal, a low-rank model was employed assuming separable spatial $U_{n}(r)$ and temporal $V_{n}(t)$ components of the metabolite signal,
@@
As final step, metabolic quantification was carried out by LCModel 47 spectral fitting, with more details mentioned in Supplementary Material.
@@ 3 Results
In Figure 2 the results of WALINET and LIPNET on simulated data are compared to L2 and HLSVD. Evaluation data were created by merging 100,000 simulated metabolite spectra with in-vivo water & lipid signals (for WALINET) or only lipid signals (for LIPNET) extracted from a subject excluded from the training data. Representative simulated evaluation spectra contaminated by water & lipid are shown in Figure 2 together with the metabolite spectra obtained by WALINET, LIPNET, and L2. Visibly, there is more agreement between ground truth metabolite spectra and the predicted metabolite spectra in the case of WALINET and LIPNET than in the case of L2 regularization. Spectra obtained by L2 regularization tend to show more residual lipid signal and more suppression of the NAA peak compared to WALINET and LIPNET. Quantitatively, boxplots of the normalized root mean square error (NRMSE) show that WALINET and LIPNET remove more lipid signal (interquartile NRMSE 0.86%-2.69%) while preserving more metabolite signal (interquartile NRMSE 0.62%-1.45%) compared to L2 (lipid interquartile NRMSE 3.68%-6.45% and metabolite interquartile NRMSE 1.04%-4.11%).
Figure 2: Simulation results. a) Comparison of lipid suppression by LIPNET and L2. Input spectra contaminated by lipids are shown in the first row (black), second row shows the metabolite spectra recovered by LIPNET (red), third row shows metabolite spectra recovered by L2 regularization (orange). b) Input spectra contaminated by water and lipids are showed on the first row (black), and the metabolite spectra predicted by WALINET (red) are plotted below. The ground-truth metabolite spectra (blue) are overlaid in all spectral plots. c) Normalized root mean squared error (NRMSE) is computed for the whole spectrum (4.7-0.7), the metabolic range (4.2-1.9ppm) and the lipid range (1.9-0.7ppm) for each method WALINET, LIPNET and L2. Separate evaluation of specific spectral ranges allows an individual assessment of the preservation of metabolic signals, as well as the effectiveness of lipid and water suppression. Arrows indicate the position of the main peaks of water, lipids and metabolites.
@@
WALINET (mm:ss) HLSVD+L2 (mm:ss) HLSVD (mm:ss) L2 (mm:ss)
@@
Metabolic maps of NAA+NAAG obtained with lipid suppression methods show similar structural features with good contrast between gray-white matter. However, some differences can be observed: 1) L2 regularization produces lower levels for NAA+NAAG maps compared to NAA+NAAG levels obtained by LIPNET, 2) the metabolic maps obtained with the strongest L2 regularization ($\beta=7.38*10^{6}$) have the lowest metabolite levels. The superior performance of LIPNET is also confirmed by the spectral quality maps, which show 10%-50% smaller CRLB and 155% higher SNR compared to L2 regularization. Lipid maps show 60% lower residual lipid signal for LIPNET compared to L2 regularization. On the other hand, it can be seen that without any lipid removal the metabolic maps are completely overwhelmed by lipid artifacts with no visible structural details of the brain and very large quantification errors. Examples of spectra show clearly more residual lipid signal by L2 regularization than by LIPNET, while in the case of no lipid removal the metabolite spectra are heavily distorted by the large lipid signal. In addition, metabolic maps of Cr+PCr, and Glutamate, which are consistent with the previous findings, are presented in Supplementary Figure 2.
Figure 3: Comparison of lipid removal on in-vivo 2D MRSI by LIPNET and L2 for three values of the regularization parameter $\beta$, the optimal value ($3.69*10^{6}$), double the optimal value, and zero for no lipid suppression. Metabolic maps are shown for NAA+NAAG, the corresponding CLRB, SNR computed by LCModel and residual lipid signal. Spectra from several brain voxels are shown for each method, the white trace shows the measured spectrum, the red trace shows LCModel fit.
Second, we studied the effects of different water removal methods. Figure 4 compares the water removal by WALINET and HLSVD on in-vivo 2D MRSI. Similar maps of the residual water signal and metabolites are obtained for WALINET and LIPNET+HLSVD. The maps obtained without water suppression show higher residual water signal and signal dropout in the center of the brain, which is worse for L2 than LIPNET. Spectra show progressively larger baseline noise and more distortion of metabolite peaks when going from WALINET to LIPNET+HLSVD, LIPNET, and L2.
@@
Figure 4: Comparison of water removal methods on in-vivo 2D MRSI, including WALINET, HLSVD+LIPNET, LIPNET and L2 without water suppression. WET11 water suppression was used during acquisition. Metabolic maps and the corresponding CRLBs are shown for NAA+NAAG, Cr+PCr and Glu together with maps of the residual water signal and examples of spectra (voxel locations indicated by arrows, red line indicates LCModel fit and the white line the experimental spectra).
@@
Figure 5: Comparison of combined water & lipid removal on in-vivo 3D MRSI. Results are shown for WALINET, HLSVD+LIPNET and HLSVD+L2, including metabolic maps for NAA, Glutamate, Inositol, residual lipid signal, residual water signal, and SNR computed by LCModel. Selected spectra from individual voxels indicated by white arrows on the anatomical images are shown at the bottom (white trace shows measured spectrum, red trace shows LCModel fit).

### v2

# WALINET: A water and lipid identification convolutional Neural Network for nuisance signal removal in 1H MR Spectroscopic Imaging

@@ Abstract
Results. WALINET is significantly faster and needs 8s for high-resolution whole-brain MRSI, compared to 42 minutes for conventional HLSVD+L2. WALINET suppresses lipid and water in the brain by 25-45 and 34-53 fold, respectively. WALINET has better performance than HLSVD+L2, providing: 1) more lipid removal with 41% lower NRMSE, 2) better metabolite signal preservation with 71% lower NRMSE in simulated data, 155% higher SNR and 50% lower CRLB in in-vivo data. Metabolic maps obtained by WALINET in healthy subjects and patients show better gray/white-matter contrast with more visible structural details.
@@ 1 Introduction
Proton magnetic resonance spectroscopic imaging (1H-MRSI) has great potential as a metabolic imaging technique that can measure the intrinsic metabolite concentrations across the whole-brain without the need to administer molecular agents bogner2021accelerated. Proton-based MR spectroscopy, compared to X-nuclei spectroscopy, has a significantly higher SNR and is ubiquitously available on all MRI scanners equipped with standard hardware and software. Consequently, it represents the vast majority of MRS performed in clinical studies and research applications. In specific pathology it allows the detection of metabolic abnormalities in the brain before anatomical lesions are visible oz2014clinical. Therefore, it is valuable for disease investigation such as brain tumor identification, classification and treatment response assessment kickingereder2018radiomics. It further enables the study of neurochemistry alterations and disease mechanisms in neuropsychiatry su2016whole. However, the clinical potential of 1H-MRSI is not fully realized due to complex experimental factors that reduce data quality such as large artifacts from the overwhelming water and lipid signals maudsley2021advanced, wilson2019methodological. Hence, efficient solutions to these technical problems can have great and widespread impact for the use of 1H-MRSI in clinical applications.
@@
The removal of lipid signal originating from the scalp region presents with even greater challenges tkavc2021water. The lipid spectrum is complicated, having multiple peaks that fully overlap with aliphatic region and are 1-2 orders of magnitude stronger than metabolite peaks. Furthermore, the lipid signal originates from scalp areas with very inhomogeneous $B_{0}$ field, hence they are much broader than water and metabolite peaks from inside the brain.
Methods enabling suppression of water and lipid contamination can be categorized into several groups tkavc2021water: techniques leveraging specific RF-pulse or sequence designs to invert, nullify or saturate water or lipid resonances balchandani2008fat, zhu2010dual, hangel2015lipid, ogg1994wet, tkavc1999vivo, esmaeili2017three, weng2022slow, chang2018non approaches utilizing dedicated hardware to spoil the scalp signal strasser2022improving, de2018elliptical, and post-acquisition methods employing spatial or spectral priors for lipid contamination removal. The latter category in the case of lipids removal can be divided in methods using lipid signal extrapolation haupt1996removal, dual-density reconstruction metzger1999hybrid, sarkar2002truncation, lipid-basis penalty bilgic2013lipid, lee2010iterative, subspace reconstruction methods based on specific spatial supports and spectral decomposition ma2016removal, adany2021method, bhattacharya2017compartmentalized, while in the case of water removal includes Hankel matrix singular value decomposition barkhuijsen1987improved, subspace reconstruction ma2016removal, Lowner tensorization nagaraja2018tensor, and water-basis penalty lin2019water.
@@
Recently, closed form solutions have been developed for lipid removal bilgic2014fast, tsai2019reduction. Closed form solutions are derived by applying a linear operator on each individual spectrum and are faster compared to iterative lipid suppression algorithms. However, the linear operator method has some drawbacks too: 1) it is subject specific, 2) requires accurate anatomical mask delineation, and 3) may not work if the orthogonal assumption and linear superposition between lipids and metabolites is not met..
@@
In recent years, deep learning-based methods applied to mitigate MRS challenges have gained popularity, enabling robust applications and eliminating the need for complex parameter optimizations. Thereby, applications can be categorized into deep learning based artifact removal, denoising, lowrank kreis2018artifactcnn, lam2019constrained, lee2019intact and spectral quantification gurbani2019incorporation, hatami2018magnetic, louis2021quantification, lee2020deep, starcuk2023quantcnn.
@@ 2 Methods
The problem can be formulated using two distinct inputs $x_{1}$ and $x_{2}$: 1) the original MRS spectrum ($x_{1}$) containing metabolite signal $m$ contaminated with lipid $l$ and water $w$ signal, and 2) the spectrum ($x_{2}$) subjected to a projection onto the lipid subspace, $\mathds{1}$-$\mathcal{L}$, using lipid L2 regularization bilgic2014fast approach with $\mathcal{L}=(\mathds{1}+\beta\mathbf{LL^{H}})^{-1}$. $\mathbf{L}$ is a matrix containing the in-vivo lipid signal obtained from the scalp mask of each subject and $\beta$ is the regularization parameter. The derived operator $\mathcal{L}$ is an approximation of a projection whose kernel is given by the linear span of $\mathbf{L}$. $\mathds{1}$-$\mathcal{L}$, with $\mathds{1}$ being an identity matrix, is a projection onto the lipid subspace span($\mathbf{L}$) (see references for further details). A separate lipid projection operator is calculated for each subject. The network inputs are defined as
@@
WALINET employs a Y-Net convolutional neural network structure mohammed2018net, depicted in Figure 1, and characterized by the use of two encoders instead of one. Note, that the same network architecture can be used and trained only for lipid removal, which we call LIPNET (LIpid neural NETwork).
The Y-Net architecture enables enhanced contextual understanding by integrating features from different branches, leading to improved understanding of complex structures within the data. The results show an improvement compared to the U-Net mohammed2018net and enable the encoding of different features in each encoder lan2020net.
@@
Each encoder ($\mathcal{E}$) and decoder ($\mathcal{D}$) of the Y-Net $(\mathcal{Y}$) comprised four convolutional blocks, each consisting of two convolutional layers, PReLU activation functions he2015delving, dropout with a rate of 0.01, and MaxPooling/upsampling with a factor of 2. Skip connections are implemented between the encoders and decoder. An additional convolutional block is incorporated in the bottleneck region and after the decoder. The outputs of the encoders are concatenated and forwarded to the bottleneck convolutional block in the decoder.
@@
The training spectra were generated through a multi-step process, where metabolite spectra were simulated and combined with experimentally measured lipid and water spectra from in-vivo human data. This approach was chosen because: 1) metabolite spectra can be realistically simulated for an extremely large range of very diverse parameters, 2) the lipid and water spectra are affected by a complex combination of experimental factors which are hard to be fully accounted in simulations, hence more realistic spectra can be extracted from measured data.
Spectra for 25 common 1H metabolites were simulated using a physical model for the coupled spin systemsSmith1994, songeon2023vivo. An extensive dataset of $1.9\times 10^{6}$ metabolite spectra was simulated for a wide range of concentrations, linewidths, noise levels, and baselines. Water and lipid signals were extracted from 19 subjects, including 2 glioma patients. The MRSI data were acquired with the 3D 1H-FID-ECCENTRIC sequence described in paragraph 2). $10^{5}$ lipid & water spectra were extracted from each subject, resulting in a total $1.9\times 10^{6}$ lipid & water spectra that were randomly combined with metabolite spectra for training. Additionally, a validation dataset included 4 other subjects. Further details regarding generation of the training data is provided in Supplementary Information.
@@
The network was trained for 400 epochs, using the Adam kingma2014adam optimizer with a learning rate of 0.01, which was quartered every 50 epochs. The exponential decay rates of the first and second momentum of Adam $\beta_{1}$ & $\beta_{2}$ were set to 0.9 and 0.999. The network was implemented using PyTorch 2.0.1 and CUDA 11.7 packages in Python 3.8.13. The model training was performed on a PowerEdge R7525 server (Dell) with 64 CPU cores (AMD EPYC7542 2.90GHz, 128M Cache, DDR4-3200), 512 GB CPU RAM (RDIMM, 3200MT/s), 3 GPU NVIDIA Ampere A40 (PCIe, 48GB GPU RAM) running Rocky Linux release 8.8 (Green Obsidian).
@@
In-vivo MRSI data were acquired with 2D 1H-FID Cartesian phase encoded Klauser2021 and 3D 1H-FID ECCENTRIC klauser2023eccentric pulse sequences using a 7T MR scanner (MAGNETOM Terra, Siemens Healthineers, Forchheim, Germany) and a 1Tx/32Rx head coil (NovaMedical, Wilmington, MA, USA).
@@
3D 1H-FID-ECCENTRIC klauser2023eccentric was acquired on 21 subjects with 64x64x31 matrix, 220x220x105 mm3 FoV, 3.4x3.4x3.4 mm3 voxel size, spectral bandwidth of 2326 Hz, and 453 FID points.
@@
[…]
@@
This value was selected as the optimal trade-off between minimizing metabolite alteration and maximizing lipid suppression.
Following the water and lipid removal, a low-rank model was employed assuming separable spatial $U_{n}(r)$ and temporal $V_{n}(t)$ components of the metabolite signal,
@@
As final step, metabolic quantification was carried out by LCModel provencher1993estimation spectral fitting.
@@ 3 Results
In Figure 2 the results of LIPNET and WALINET on simulated data are compared to L2 and HLSVD+L2, respectively. Evaluation data were created by merging 100,000 simulated metabolite spectra with in-vivo water & lipid signals (for WALINET) or only lipid signals (for LIPNET) extracted from a subject excluded from the training data. Results on evaluation spectra contaminated only by lipid are shown in Figure 2a, and by water & lipid are shown in Figure 2b. We obtained more agreement between ground truth metabolite spectra and the predicted metabolite spectra in the case of LIPNET and WALINET than in the case of L2 and HLSVD+L2 regularization. Spectra obtained by L2 regularization tend to show more residual lipid signal and more suppression of the NAA peak compared to WALINET and LIPNET. Quantitatively, boxplots of the normalized root mean square error (NRMSE) show that WALINET and LIPNET remove more lipid signal (interquartile NRMSE 0.86%-2.69%) while preserving more metabolite signal (interquartile NRMSE 0.62%-1.45%) compared to L2 (lipid interquartile NRMSE 3.68%-6.45% and metabolite interquartile NRMSE 1.04%-4.11%).
Figure 2: Simulation results. a) Comparison of lipid suppression by LIPNET and L2. Input spectra contaminated by lipids are shown in the first row (black), second row shows metabolite spectra recovered by L2 regularization (orange), third row shows the metabolite spectra recovered by LIPNET (red). b) Input spectra contaminated by water and lipids are shown on the first row (black), in the second row metabolite spectra processed with HLSVD+L2 are displayed, and the metabolite spectra predicted by WALINET (red) are plotted below. The ground-truth metabolite spectra (blue) are overlaid in all spectral plots. Normalized root mean squared error (NRMSE) is computed for the whole spectrum (9.0-0.0 ppm), the metabolic range (4.2-1.9 ppm) and the lipid range (1.9-0.7 ppm) for each method LIPNET or L2 in a) and WALINET or HLSVD+L2 in b). Separate evaluation of specific spectral ranges allows an individual assessment of the preservation of metabolic signals, as well as the effectiveness of lipid and water suppression. Arrows indicate the position of the main peaks of water, lipids and metabolites.
@@
WALINET
($m$:$s$)
HLSVD+L2
($m$:$s$)
HLSVD
($m$:$s$)
L2
($m$:$s$)
@@
Metabolic maps of NAA+NAAG obtained with lipid suppression methods show similar structural features with good contrast between gray-white matter. However, some differences can be observed: 1) L2 regularization produces lower levels for NAA+NAAG maps compared to NAA+NAAG levels obtained by LIPNET, 2) the metabolic maps obtained with the strongest L2 regularization ($\beta=7.38*10^{6}$) have the lowest metabolite levels. The superior performance of LIPNET is also confirmed by the spectral quality maps, which show 10%-50% smaller CRLB and 155% higher SNR compared to L2 regularization. Lipid maps show 60% lower residual lipid signal for LIPNET compared to L2 regularization. On the other hand, it can be seen that without any lipid removal the metabolic maps are completely overwhelmed by lipid artifacts with no visible structural details of the brain and very large quantification errors. The lipid Examples of spectra show clearly more residual lipid signal by L2 regularization than by LIPNET, while in the case of no lipid removal the metabolite spectra are heavily distorted by the large lipid signal. In addition, metabolic maps of Cr+PCr, and Glutamate, which are consistent with the previous findings, are presented in Supplementary Figure S2.
Figure 3: Comparison of lipid removal on in-vivo 2D MRSI by LIPNET and L2 for three values of the regularization parameter $\beta$, the optimal value ($3.69*10^{6}$), double the optimal value, and zero for no lipid suppression. Metabolic maps are shown for NAA+NAAG, the corresponding CLRB, SNR computed by LCModel provencher1993estimation and residual lipid signal. Spectra from several brain voxels are shown for each method, the white trace shows the measured spectrum, the red trace shows LCModel fit.
Second, we studied the effects of different water removal methods. Figure 4 compares the water removal by WALINET and HLSVD on in-vivo 2D MRSI. Similar maps of the residual water signal and metabolites are obtained for WALINET and LIPNET+HLSVD. The maps obtained without water suppression show higher residual water signal and signal dropout in the center of the brain, which is worse for L2 than LIPNET. Spectra and LCModel provencher1993estimation fit show progressively larger baseline noise and more distortion of metabolite peaks when going from WALINET to LIPNET+HLSVD, LIPNET, and L2. Examples of spectra with full spectral window and intensity range are shown in Supplementary Figure S3.
@@
Figure 4: Comparison of water removal methods on in-vivo 2D MRSI, including WALINET, HLSVD+LIPNET, LIPNET and L2 without water suppression. WETogg1994wet water suppression was used during acquisition. Metabolic maps and the corresponding CRLBs are shown for NAA+NAAG, Cr+PCr and Glu together with maps of the residual water signal and examples of spectra (voxel locations indicated by arrows, red line indicates LCModel provencher1993estimation fit and the white line the experimental spectra).
@@
Figure 5: Comparison of combined water & lipid removal on in-vivo 3D MRSI. Results are shown for WALINET, HLSVD+LIPNET and HLSVD+L2, including metabolic maps for NAA, Glutamate, Inositol, residual lipid signal, residual water signal, and SNR computed by LCModel. Selected spectra from individual voxels indicated by white arrows on the anatomical images are shown at the bottom (white trace shows measured spectrum, red trace shows LCModel provencher1993estimation fit).
@@
The lipid removal factor and water removal factor was quantified in-vivo and simulations for each of the methods and listed in Supplementary Table S1. It can be seen that WALINET and LIPNET provide in-vivo a mean lipid removal factor between 25-45 in the brain (689-876 lipid removal factor in the scalp), and a mean water removal factor between 34-53.